Ticket Number: HEA-NOC/20080408-6 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080408 17:31 UTC+1 Problem Start: 20080410 09:00 UTC+1
Ticket Update: 20080409 17:56 UTC+1 Problem End: 20080410 12:00 UTC+1
Our connection to INEX Lan #2 will move to our new IP equipment on Thursday
Traffic to some INEX peers will reroute via INEX Lan #1 (our connection to
which is in Citywest) or, as a last resort, via the general internet.
20080408 17:29 DW Open ticket. Contingent on our connection clearing
quarantine tomorrow, we will preconfigure our router tomorrow, and shut
down the old interface on Thursday morning prior to bringing up the new
one and transferring the BGP sessions.
20080409 17:56 DW Cleared quarantine, can ping other INEX routers. Our
router has been preconfigured, ready to proceed tomorrow morning.
Time to Fix:
3 hours scheduled.
This ticket can be monitored at http://www.hea.net/tickets/20080408-6
HEAnet Limited. Registered in Ireland, No. 275301.